caculate the linear expansivity of brass of length 120m, that assume a new length of 120.05m when headed through a temperature of 100


1
Expert's answer
2022-01-28T08:14:36-0500

*Assuming the temperature change is 100°C.

"\alpha=\dfrac{L-L_0}{L_0\Delta T}=4.167·10^{-16}\text{ 1/K}."